A 1100 kg racing car accelerates from rest at a constant rate and covers a distance of 675 m in 15 s. What is the car's acceleration? (in m/s2)
Answer #1

initial velocity = V0 = 0 m/sec ( rest)

distance = S = 675 m
time = t = 15 sec

             S = V0 *t + 1/2 * a t^2

              675 = 0 + 1/2 * a * 15^2

               675 = 112.5 *a

acceleration = a = 675 / 112.5 = 6 m/sec^2